# Quantitative Propagation of Chaos for 2D Viscous Vortex Model with General Circulations on the Whole Space

### v1

@@ 2.1. Time Evolution of Relative Entropy
$\displaystyle$== $\displaystyle-\frac{\sigma}{k}\sum_{i=1}^{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G^{N,k}\Big{|}\nabla_{x_{i}}\log\frac{G^{N,k}}{g^{k}}\Big{|}^{2}\,\mathrm{d}Z^{k}$
$\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\frac{1}{N}\sum_{j=1}^{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}G^{N,k}\Big{(}m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{k}$
$\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k+1}}G^{N,k+1}\Big{(}m_{k+1}K(x_{i}-x_{k+1})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{k+1}$
$\displaystyle+\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k+1}}G^{N,k+1}\Big{(}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}\log G^{N,k}\Big{)}\,\mathrm{d}Z^{k+1}.$
@@
$+\frac{N-k}{N}\sum_{i=1}^{k}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}G^{N,k+1}\,\mathrm{d}z_{k+1}=\sum_{i=1}^{k}\Delta_{x_{i}}G^{N,k}.$
while the tensorized law $g^{k}$ solves (1.8):
@@
$\displaystyle$=\,= $\displaystyle\frac{1}{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}\partial_{t}G^{N,k}(\log G^{N,k}+1-\log g^{k})+\frac{G^{N,k}}{g^{k}}\partial_{t}g^{k}\,\mathrm{d}Z^{k}$
$\displaystyle$=\,= $\displaystyle\frac{1}{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}\Big{(}\sigma\sum_{i=1}^{k}\Delta_{x_{i}}G^{N,k}-\frac{1}{N}\sum_{i,j=1}^{k}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}G^{N,k}$
$\displaystyle\qquad-\frac{N-k}{N}\sum_{i=1}^{k}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}G^{N,k+1}\,\mathrm{d}z_{k+1}\Big{)}(\log G^{N,k}+1-\log g^{k})\,\mathrm{d}Z^{k}$
$\displaystyle+\frac{1}{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}\Big{(}\sigma\sum_{i=1}^{k}\Delta_{x_{i}}g^{k}-\sum_{i=1}^{k}K\ast\omega(x_{i})\cdot\nabla_{x_{i}}g^{k}\Big{)}\frac{G^{N,k}}{g^{k}}\,\mathrm{d}Z^{k}.$
@@
$\displaystyle-\frac{\sigma}{k}\sum_{i=1}^{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}\nabla_{x_{i}}G^{N,k}\cdot\nabla_{x_{i}}(\log G^{N,k}-\log g^{k})\,\mathrm{d}Z^{k}$
$\displaystyle+\frac{1}{k}\sum_{i=1}^{k}\frac{1}{N}\sum_{j=1}^{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}(\log G^{N,k}-\log g^{k})G^{N,k}\,\mathrm{d}Z^{k}$
$\displaystyle+\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k+1}}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}(\log G^{N,k}-\log g^{k})G^{N,k+1}\,\mathrm{d}Z^{k+1}$
$\displaystyle-\frac{\sigma}{k}\sum_{i=1}^{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}\nabla_{x_{i}}g^{k}\cdot\nabla_{x_{i}}\frac{G^{N,k}}{g^{k}}\,\mathrm{d}Z^{k}$
$\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{k}}K\ast\omega(x_{i})\cdot\nabla_{x_{i}}g^{k}G^{N,k}\,\mathrm{d}Z^{k}.$
@@ 2.2. Law of Large Numbers and Large Deviation Type Theorems
In the original paper [JW18], Jabin-Wang introduced two Law of Large Numbers and Large Deviation type theorems, which are crucial in controlling the expectation of some mean-zero functions with respect to the joint law, bounding it with the normalized relative entropy itself and some $O(1/N)$ error terms, while one may only get $O(1)$ bound by crude estimates. Their method includes expanding the exponential term by Taylor’s expansion and counting carefully by combinatory techniques that how many terms actually do not vanish, using the one-side or two-side cancellation rules. Note that in our setting, since we should view $(m,x)$ together as a single variable $z$ to keep the symmetry, the two theorems should be changed formally, see also [SZ24].
@@
Consider any probability density $\bar{\rho}(z)$ on ×${\mathbb{R}}\times{\mathbb{R}}^{2}$ and any test function $\phi(z,w)\in L^{\infty}(({\mathbb{R}}\times{\mathbb{R}}^{2})^{2})$ satisfying the one-side cancelling property
×$\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}\phi(z,w)\bar{\rho}(w)\,\mathrm{d}w=0,\,\mbox{for any\, }z,$ (2.1)
@@
(×$\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\bar{\rho}^{\otimes N}\exp\Big{(}\frac{1}{N}\sum_{j,k=1}^{N}\phi(z_{1},z_{j})\phi(z_{1},z_{k})\Big{)}\,\mathrm{d}Z^{N}\leq C<\infty.$
@@
Consider any probability density $\bar{\rho}(z)$ on ×${\mathbb{R}}\times{\mathbb{R}}^{2}$ and any test function $\phi(z,w)$ on (×$({\mathbb{R}}\times{\mathbb{R}}^{2})^{2}$ satisfying the two-side cancelling properties
@@
where the constant $C_{JW}=1600^{2}+36e^{4}$ is obtained in [JW18, Theorem 4]. Then we have
@@
We refer the proofs of the above theorems to [JW18].
The Law of Large Numbers and Large Deviation Theorems above have been successfully applied to estimate the global relative entropy on the torus for various models. But in the setting of this paper, there are two main difficulties to be overcome.
The first problem is that we are now working on the whole space case, hence we can no longer expect that the test function $\phi$ in Theorem 2.1 is bounded. We also mention that Lim-Lu-Nolen [LLN20] provides another proof for Theorem 2.2 using martingale inequalities when $\phi\in L^{\infty}$. In the recent work Du-Li [DL24] dealing with a new collision-oriented particle system for approximating the Landau type equations, the Law of Large Numbers Theorem 2.1 is extended to the case that $\phi$ has linear growth in the second variable $w$ uniformly in the first variable $z$, using the classical Hoeffding’s inequality in high-dimensional probability, see for instance the classic book by Vershynin [Ver18]. A special case of this extension is also proved in the recent work of Gong-Wang-Xie [GWX24]. This result should be crucial to our method.
The second problem is that in order to obtain the local optimal rate $O(k^{2}/N^{2})$ for the relative entropy estimate, we need to scale the small constant $\gamma$ by a factor of $k/N$, see Subsection 2.4 below. However, the upper bound for the functional firstly obtained in [JW18] can not well reflect this scaling, hence we need to estimate the upper bound more carefully. This has been observed by S. Wang [Wan24] for Theorem 2.2, and we are now giving the similar improvements for both theorems.
@@
Consider any probability density $\bar{\rho}(z)$ on ×${\mathbb{R}}\times{\mathbb{R}}^{2}$ and any test function $\phi(z,w)$ on (×$({\mathbb{R}}\times{\mathbb{R}}^{2})^{2}$ vanishing on the diagonal $\Delta=\{(z,z):z\in{\mathbb{R}}\times{\mathbb{R}}^{2}\}$ and satisfying the one-side cancelling property
×$\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}\phi(z,w)\bar{\rho}(w)\,\mathrm{d}w=0,\,\mbox{for any\, }z,$ (2.4)
@@
$\inf\Big{\{}c>0:\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}\exp\Big{(}\sup_{z}|\phi(z,w)|^{2}/c^{2}\Big{)}g(w)\,\mathrm{d}w\Big{\}}<\lambda c_{0},$
@@
$\log\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\bar{\rho}^{\otimes N}\exp\Big{(}\frac{1}{N}\sum_{j,k=1}^{N}\phi(z_{1},z_{j})\phi(z_{1},z_{k})\Big{)}\,\mathrm{d}Z^{N}\leq\frac{C}{\lambda}<\infty.$
@@
Consider any probability density $\bar{\rho}(z)$ on ×${\mathbb{R}}\times{\mathbb{R}}^{2}$ and any test function $\phi(z,w)$ on (×$({\mathbb{R}}\times{\mathbb{R}}^{2})^{2}$ vanishing on the diagonal $\Delta=\{(z,z):z\in{\mathbb{R}}\times{\mathbb{R}}^{2}\}$ and satisfying the two-side cancelling properties
@@
where the constant $C_{JW}=1600^{2}+36e^{4}$ is obtained in [JW18, Theorem 4]. Then we have
@@
The proof of the extended Law of Large Numbers Theorem 2.3 is referred to [DL24], where the constant upper bound is slightly modified, while the proof of the extended Large Deviation Theorem 2.4 is referred to [Wan24] without any changes.
Finally we present two lemmas, say the Gibbs inequality for changing the probability measure for integration which is closely related to the variational expression for the relative entropy, and the chain rule of the relative entropy for disintegrated probability measures.
Lemma 2.3 (Gibbs Inequality).
Consider any probability density $\bar{\rho}(z)$ on ×${\mathbb{R}}\times{\mathbb{R}}^{2}$ and $\rho_{N}(Z^{N})$ on (×$({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}$ and any test function $\Phi(Z^{N})$ on (×$({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}$, then for any $\eta>0$ we have
(×$\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Phi\rho_{N}\,\mathrm{d}Z^{N}\leq\frac{1}{\eta}H_{N}(\rho_{N}|\bar{\rho}^{N})+\frac{1}{\eta N}\log\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\exp(N\eta\Phi)\,\mathrm{d}Z^{N}.$
Lemma 2.4 (Chain Rule, borrowed from [Lac23]).
@@
We refer the proof of Lemma 2.3 to [JW18] and the proof of Lemma 2.4 to Budhiraja-Dupuis [BD19].
@@ 2.3. Global Relative Entropy Control
In this subsection, with all the key tools for preparations in the last subsection, we shall prove our main estimate in the right-handside of Lemma 2.1. Thanks to the non-positive dissipation term, we only need to control the second term, which we recall
$-\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{(}m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{N}.$
Due to the appearance of $m_{j}$ in front of the singular kernel $K(x_{i}-x_{j})$, it seems impossible to use the symmetrization trick as in [JW18, FW23]. Hence we use integration by parts, thanks to that $K\in W^{-1,\infty}$, with $K=\nabla\cdot V$ as in [JW18] where
@@
$\displaystyle\,\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\Big{(}\nabla_{x_{i}}g^{N}\otimes\nabla_{x_{i}}\frac{G^{N}}{g^{N}}\Big{)}\,\mathrm{d}Z^{N}$ (2.7)
$\displaystyle$++ $\displaystyle\,\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\frac{\nabla_{x_{i}}^{2}g^{N}}{g^{N}}\,\mathrm{d}Z^{N}.$ (2.8)
@@
It is direct to check that $\phi$ satisfies the two cancelling conditions (2.2). Also we have $\phi$ satisfy the growth condition
$f(m,x)\triangleq\sup_{m^{\prime},x^{\prime}}|\phi(m,x,m^{\prime},x^{\prime})|\leq C\Big{(}\frac{1+\log(1+t)}{1+t}+\frac{|x|^{2}}{(1+t)^{2}}\Big{)}$
since $V$ is bounded and ${\mathcal{M}}$ is compactly supported, and $\nabla^{2}g/g$ satisfies the quadratic growth estimate as in Proposition 4.1 below. Moreover, for any $\lambda(t)>0$, as in [JW16] we have
×$\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}e^{\lambda(t)f}g(z)\,\mathrm{d}z\geq\frac{\lambda^{p}}{p!}\|f\|_{L^{p}(g\,\mathrm{d}z)}^{p}\geq\frac{\lambda^{p}}{p^{p}}\|f\|_{L^{p}(g\,\mathrm{d}z)}^{p}.$
Hence by choosing $\lambda(t)=\varepsilon(1+t)/(1+\log(1+t))$ with $\varepsilon>0$ small enough we have
$\displaystyle\sup_{p\geq 1}\frac{\|f\|_{L^{p}(g\,\mathrm{d}z)}}{p}\leq\frac{1}{\lambda}\int_{{\mathbb{R}}\times{\mathbb{R}}^{2}}e^{\lambda f}g(z)\,\mathrm{d}z$
@@
where $C_{\varepsilon}$ is some large constant depending on $\varepsilon$, and we use the Gaussian upper bound of $g$ as in Lemma 3.2 below. Applying the Gibbs inequality Lemma 2.3 for the test function
@@
and for some $\eta(t)=C_{\varepsilon}^{-1}\frac{1+t}{1+\log(1+t)}$ and using the Large Deviation Theorem 2.2, we bound the second term (2.8) after fixing some small $\varepsilon$ by
$C\frac{1+\log(1+t)}{1+t}\Big{(}H_{N}(G^{N}|g^{N})+\frac{1}{N}\Big{)}.$
Control of (2.7) : For the first term (2.7), we first use a Cauchy-Schwartz argument to bound with the normalized Fisher information and some error term
$\displaystyle\frac{\sigma}{4N}\sum_{i=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{|}\nabla_{x_{i}}\log\frac{G^{N}}{g^{N}}\Big{|}^{2}\,\mathrm{d}Z^{N}$ (2.9)
$\displaystyle$++ $\displaystyle\frac{2}{N\sigma}\sum_{i=1}^{N}\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{|}\frac{1}{N}\sum_{j=1}^{N}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}\Big{|}^{2}|\nabla\log g(m_{i},x_{i})|^{2}\,\mathrm{d}Z^{N}.$ (2.10)
The first term (2.9) is cancelled by the dissipation term in Lemma 2.1. Hence we only need to control the second term (2.10). Due to symmetry, we rewrite it as
(×$\int_{({\mathbb{R}}\times{\mathbb{R}}^{2})^{N}}G^{N}\Big{|}\frac{1}{N}\sum_{j=1}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}.$

### v2

@@ 2.1. Time Evolution of Relative Entropy
$\displaystyle$== $\displaystyle-\frac{\sigma}{k}\sum_{i=1}^{k}\int_{\mathbb{D}^{k}}G^{N,k}\Big{|}\nabla_{x_{i}}\log\frac{G^{N,k}}{g^{k}}\Big{|}^{2}\,\mathrm{d}Z^{k}$
$\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\frac{1}{N}\sum_{j=1}^{k}\int_{\mathbb{D}^{k}}G^{N,k}\Big{(}m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{k}$
$\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{\mathbb{D}^{k+1}}G^{N,k+1}\Big{(}m_{k+1}K(x_{i}-x_{k+1})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{k+1}$
$\displaystyle+\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{\mathbb{D}^{k+1}}G^{N,k+1}\Big{(}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}\log G^{N,k}\Big{)}\,\mathrm{d}Z^{k+1}.$
@@
$+\frac{N-k}{N}\sum_{i=1}^{k}\int_{\mathbb{D}}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}G^{N,k+1}\,\mathrm{d}z_{k+1}=\sum_{i=1}^{k}\Delta_{x_{i}}G^{N,k}.$
while the factorized law $g^{k}$ solves (1.8):
@@
$\displaystyle$=\,= $\displaystyle\frac{1}{k}\int_{\mathbb{D}^{k}}\partial_{t}G^{N,k}(\log G^{N,k}+1-\log g^{k})+\frac{G^{N,k}}{g^{k}}\partial_{t}g^{k}\,\mathrm{d}Z^{k}$
$\displaystyle$=\,= $\displaystyle\frac{1}{k}\int_{\mathbb{D}^{k}}\Big{(}\sigma\sum_{i=1}^{k}\Delta_{x_{i}}G^{N,k}-\frac{1}{N}\sum_{i,j=1}^{k}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}G^{N,k}$
$\displaystyle\qquad-\frac{N-k}{N}\sum_{i=1}^{k}\int_{\mathbb{D}}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}G^{N,k+1}\,\mathrm{d}z_{k+1}\Big{)}(\log G^{N,k}+1-\log g^{k})\,\mathrm{d}Z^{k}$
$\displaystyle+\frac{1}{k}\int_{\mathbb{D}^{k}}\Big{(}\sigma\sum_{i=1}^{k}\Delta_{x_{i}}g^{k}-\sum_{i=1}^{k}K\ast\omega(x_{i})\cdot\nabla_{x_{i}}g^{k}\Big{)}\frac{G^{N,k}}{g^{k}}\,\mathrm{d}Z^{k}.$
@@
$\displaystyle-\frac{\sigma}{k}\sum_{i=1}^{k}\int_{\mathbb{D}^{k}}\nabla_{x_{i}}G^{N,k}\cdot\nabla_{x_{i}}(\log G^{N,k}-\log g^{k})\,\mathrm{d}Z^{k}$
$\displaystyle+\frac{1}{k}\sum_{i=1}^{k}\frac{1}{N}\sum_{j=1}^{k}\int_{\mathbb{D}^{k}}m_{j}K(x_{i}-x_{j})\cdot\nabla_{x_{i}}(\log G^{N,k}-\log g^{k})G^{N,k}\,\mathrm{d}Z^{k}$
$\displaystyle+\frac{1}{k}\sum_{i=1}^{k}\frac{N-k}{N}\int_{\mathbb{D}^{k+1}}m_{k+1}K(x_{i}-x_{k+1})\cdot\nabla_{x_{i}}(\log G^{N,k}-\log g^{k})G^{N,k+1}\,\mathrm{d}Z^{k+1}$
$\displaystyle-\frac{\sigma}{k}\sum_{i=1}^{k}\int_{\mathbb{D}^{k}}\nabla_{x_{i}}g^{k}\cdot\nabla_{x_{i}}\frac{G^{N,k}}{g^{k}}\,\mathrm{d}Z^{k}$
$\displaystyle-\frac{1}{k}\sum_{i=1}^{k}\int_{\mathbb{D}^{k}}K\ast\omega(x_{i})\cdot\nabla_{x_{i}}g^{k}G^{N,k}\,\mathrm{d}Z^{k}.$
@@ 2.2. Law of Large Numbers and Large Deviation Type Theorems
In the original paper [30], Jabin-Wang introduced two Law of Large Numbers and Large Deviation type theorems, which are crucial in controlling the expectation of some mean-zero functions with respect to the joint law, bounding it with the normalized relative entropy itself and some $O(1/N)$ error terms, while one may only get $O(1)$ bound by crude estimates. Their method includes expanding the exponential term by Taylor’s expansion and counting carefully by combinatorial techniques that how many terms actually do not vanish, using the one-side or two-side cancellation rules. Note that in our setting, since we should view $(m,x)$ together as a single variable $z$ to keep the symmetry, the two theorems should be changed formally, see also [52].
@@
Consider any probability density $\bar{\rho}(z)$ on $\mathbb{D}$ and any test function $\phi(z,w)\in L^{\infty}(\mathbb{D}^{2})$ satisfying the one-side cancelling property
$\int_{\mathbb{D}}\phi(z,w)\bar{\rho}(w)\,\mathrm{d}w=0,\,\mbox{for any\, }z,$ (2.1)
@@
$\int_{\mathbb{D}^{N}}\bar{\rho}^{\otimes N}\exp\Big{(}\frac{1}{N}\sum_{j,k=1}^{N}\phi(z_{1},z_{j})\phi(z_{1},z_{k})\Big{)}\,\mathrm{d}Z^{N}\leq C<\infty.$
@@
Consider any probability density $\bar{\rho}(z)$ on $\mathbb{D}$ and any test function $\phi(z,w)$ on $\mathbb{D}^{2}$ satisfying the two-side cancelling properties
@@
where the constant $C_{JW}=1600^{2}+36e^{4}$ is obtained in [30, Theorem 4]. Then we have
@@
We refer the proofs of the above theorems to [30].
The Law of Large Numbers and Large Deviation Theorems above have been successfully applied to estimate the global relative entropy on the torus for various models. But in the setting of this article, there are two main difficulties to overcome.
The first problem is that we are now working on the whole space, hence we can no longer expect that the test function $\phi$ in Theorem 2.1 is bounded. We also mention that Lim-Lu-Nolen [38] provides another proof for Theorem 2.2 using martingale inequalities when $\phi\in L^{\infty}$. In the recent work Du-Li [15] dealing with a new collision-oriented particle system for approximating the Landau type equations, the Law of Large Numbers Theorem 2.1 is extended to the case that $\phi$ has linear growth in the second variable $w$ uniformly in the first variable $z$, using the classical Hoeffding’s inequality in high-dimensional probability, see for instance the classic book by Vershynin [54]. A special case of this extension is also proved in the recent work of Gong-Wang-Xie [19]. This result should be crucial to our method.
The second problem is that in order to obtain the local optimal rate $O(k^{2}/N^{2})$ for the relative entropy estimate, we need to scale the small constant $\gamma$ by a factor of $k/N$, see Subsection 2.4 below. However, the upper bound for the functional first obtained in [30] can not well reflect this scaling, so we need to estimate the upper bound more carefully. This has been observed by S. Wang [55] for Theorem 2.2, and we are now giving similar improvements for both theorems.
@@
Consider any probability density $\bar{\rho}(z)$ on $\mathbb{D}$ and any test function $\phi(z,w)$ on $\mathbb{D}^{2}$ vanishing on the diagonal $\Delta=\{(z,z):z\in\mathbb{D}\}$ and satisfying the one-side cancelling property
$\int_{\mathbb{D}}\phi(z,w)\bar{\rho}(w)\,\mathrm{d}w=0,\,\mbox{for any\, }z,$ (2.4)
@@
$\inf\Big{\{}c>0:\int_{\mathbb{D}}\exp\Big{(}\sup_{z}|\phi(z,w)|^{2}/c^{2}\Big{)}g(w)\,\mathrm{d}w\Big{\}}<\lambda c_{0},$
@@
$\log\int_{\mathbb{D}^{N}}\bar{\rho}^{\otimes N}\exp\Big{(}\frac{1}{N}\sum_{j,k=1}^{N}\phi(z_{1},z_{j})\phi(z_{1},z_{k})\Big{)}\,\mathrm{d}Z^{N}\leq\frac{C}{\lambda}<\infty.$
@@
Consider any probability density $\bar{\rho}(z)$ on $\mathbb{D}$ and any test function $\phi(z,w)$ on $\mathbb{D}^{2}$ vanishing on the diagonal $\Delta=\{(z,z):z\in\mathbb{D}\}$ and satisfying the two-side cancelling properties
@@
where the constant $C_{JW}=1600^{2}+36e^{4}$ is obtained in [30, Theorem 4]. Then we have
@@
The proof of the extended Law of Large Numbers Theorem 2.3 is referred to [15], where the constant upper bound is slightly modified, while the proof of the extended Large Deviation Theorem 2.4 is referred to [55] without any changes.
Finally we present two lemmas, say the Donsker-Varadhan inequality for the variational expression for the relative entropy, and the chain rule of the relative entropy for disintegrated probability measures.
Lemma 2.3 (Donsker-Varadhan Inequality).
Consider any probability density $\bar{\rho}(z)$ on $\mathbb{D}$ and $\rho_{N}(Z^{N})$ on $\mathbb{D}^{N}$ and any test function $\Phi(Z^{N})$ on $\mathbb{D}^{N}$, then for any $\eta>0$ we have
$\int_{\mathbb{D}^{N}}\Phi\rho_{N}\,\mathrm{d}Z^{N}\leq\frac{1}{\eta}H_{N}(\rho_{N}|\bar{\rho}^{N})+\frac{1}{\eta N}\log\int_{\mathbb{D}^{N}}\exp(N\eta\Phi)\,\mathrm{d}Z^{N}.$
Lemma 2.4 (Chain Rule, borrowed from [34]).
@@
We refer the proof of Lemma 2.3 to [30] and the proof of Lemma 2.4 to Budhiraja-Dupuis [8].
@@ 2.3. Global Relative Entropy Control
In this subsection, with all the key tools for preparation in the last subsection, we shall prove our main estimate on the right-hand side of Lemma 2.1. Thanks to the non-positive dissipation term, we only need to control the second term, which we recall
$-\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{(}m_{j}K(x_{i}-x_{j})-K\ast\omega(x_{i})\Big{)}\cdot\nabla\log g(m_{i},x_{i})\,\mathrm{d}Z^{N}.$
Due to the appearance of $m_{j}$ in front of the singular kernel $K(x_{i}-x_{j})$, it seems impossible to use the symmetrization trick as in [30, 16]. Hence we use integration by parts, thanks to that $K\in W^{-1,\infty}$, with $K=\nabla\cdot V$ as in [30] where
@@
$\displaystyle\,\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\Big{(}\nabla_{x_{i}}g^{N}\otimes\nabla_{x_{i}}\frac{G^{N}}{g^{N}}\Big{)}\,\mathrm{d}Z^{N}$ (2.7)
$\displaystyle$++ $\displaystyle\,\frac{1}{N^{2}}\sum_{i,j=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}:\frac{\nabla_{x_{i}}^{2}g^{N}}{g^{N}}\,\mathrm{d}Z^{N}.$ (2.8)
@@
It is straightforward to check that $\phi$ satisfies the two cancelling properties (2.2). Also, we have $\phi$ satisfies the growth condition
$f(m,x)\triangleq\sup_{m^{\prime},x^{\prime}}|\phi(m,x,m^{\prime},x^{\prime})|\leq C\frac{B_{1}}{1+t}(B_{2}-\log(1+t)-\log g(m,x))$
since $V$ is bounded and ${\mathcal{M}}$ is compactly supported and $\nabla^{2}g/g$ satisfies the logarithmic estimates as in Proposition 4.1 below. Moreover, for any $\lambda(t)>0$, as in [28] we have
$\int_{\mathbb{D}}e^{\lambda(t)f}g(z)\,\mathrm{d}z\geq\frac{\lambda^{p}}{p!}\|f\|_{L^{p}(g\,\mathrm{d}z)}^{p}\geq\frac{\lambda^{p}}{p^{p}}\|f\|_{L^{p}(g\,\mathrm{d}z)}^{p}.$
Hence, in order to obtain the convergence of the exponential integral, we choose
$\lambda(t)=\frac{1+t}{2CB_{1}}$
to conclude that
$\displaystyle\sup_{p\geq 1}\frac{\|f\|_{L^{p}(g\,\mathrm{d}z)}}{p}\leq$ $\displaystyle\,\frac{1}{\lambda}\int_{\mathbb{D}}e^{\lambda f}g(z)\,\mathrm{d}z$
@@
where we use the Gaussian upper bound of $g$ as in Lemma 3.2 below. Applying the Donsker-Varadhan inequality in Lemma 2.3 for the test function
@@
and for some $\eta(t)=C^{-1}(1+t)$, and using the Large Deviation Theorem 2.4, we bound the second term (2.8) by
$\frac{C}{1+t}\Big{(}H_{N}(G^{N}|g^{N})+\frac{1}{N}\Big{)}.$
Control of (2.7) : For the first term (2.7), we first use a Cauchy-Schwarz argument to bound with the normalized relative Fisher information and some error term:
$\displaystyle\frac{\sigma}{4N}\sum_{i=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{|}\nabla_{x_{i}}\log\frac{G^{N}}{g^{N}}\Big{|}^{2}\,\mathrm{d}Z^{N}$ (2.9)
$\displaystyle$++ $\displaystyle\frac{2}{N\sigma}\sum_{i=1}^{N}\int_{\mathbb{D}^{N}}G^{N}\Big{|}\frac{1}{N}\sum_{j=1}^{N}\Big{(}m_{j}V(x_{i}-x_{j})-V\ast\omega(x_{i})\Big{)}\Big{|}^{2}|\nabla\log g(m_{i},x_{i})|^{2}\,\mathrm{d}Z^{N}.$ (2.10)
The first term (2.9) is cancelled by the dissipation term in Lemma 2.1. Therefore, we only need to control the second term (2.10). Due to symmetry, we rewrite it into
$\int_{\mathbb{D}^{N}}G^{N}\Big{|}\frac{1}{N}\sum_{j=1}^{N}\Big{(}m_{j}V(x_{1}-x_{j})-V\ast\omega(x_{1})\Big{)}\Big{|}^{2}|\nabla\log g(m_{1},x_{1})|^{2}\,\mathrm{d}Z^{N}.$